OpenStudy (♥ashley♥):

What is the next term in the pattern? 15, 16, 18, 21, 25, …

OpenStudy (mathmale):

Adding what number to 15 results in what number? Additing what number to 16 results in what number? ... and so on. Writing out this pattern will likely show you how to answer this question.

OpenStudy (rcp031):

Check out the difference between adjacent 16-15, 18-16, 21-18, 25-21 it follows a pattern 1,2,3,4 ,...... proceeds further

OpenStudy (♥ashley♥):

Is it 40

OpenStudy (rcp031):

15+1=16, 16+2=18, 18+3=21, 21+4=25, do next

OpenStudy (♥ashley♥):

25+5= 30. I meant 30 not 40

jhonyy9 (jhonyy9):

yes 30 is right sure